OpenStudy (anonymous):

what is the estimate for 7 4/9 *63?

OpenStudy (whpalmer4):

THE estimate? Wasn't aware there was only one possible estimate :-) Were I to estimate that, I would probably look at 7 4/9 and call it 7 1/2 (or 7.5), and then depending on the accuracy desired, look at the 63 and call it 60. 7.5*60 = ?

OpenStudy (mathstudent55):

7 4/9 is close to 7.5 7 x 60 = 420 8 x 60 = 480 My estimate is 460

OpenStudy (whpalmer4):

exact answer is 469. 7.5*60 = 450, and you can easily improve that a bit by adding 3*7.5 = 22.5 which gives you 472.5.

OpenStudy (whpalmer4):

the last part is brought to you by 7.5*63 = 7.5(60 + 3) = 7.5*60 + 7.5*3.

OpenStudy (anonymous):

thx